OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!! Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4) PLEASE GIVE THE EXACT ANSWER

OpenStudy (mertsj):

Can you find the slope?

OpenStudy (ybarrap):

Find the slope m and use one of the points, say <7,2>, and solve y=mx + b for the y-intercept, b: $$ 2=\cfrac{4-2}{5-7}\times7+b\\ b=2-\cfrac{4-2}{5-7}\times7\\ $$ Where $$ m=\cfrac{4-2}{5-7} $$ Once you have b and m put into the following equation: $$ y=mx+b $$